### v1

# Graph Representation Learning via Causal Diffusion for Out-of-Distribution Recommendation

Abstract
Graph Neural Networks (GNNs)-based recommendation algorithms typically assume that training and testing data are drawn from independent and identically distributed (IID) spaces. However, this assumption often fails in the presence of out-of-distribution (OOD) data, resulting in significant performance degradation. In this study, we construct a Structural Causal Model (SCM) to analyze interaction data, revealing that environmental confounders (e.g., the COVID-19 pandemic) lead to unstable correlations in GNN-based models, thus impairing their generalization to OOD data. To address this issue, we propose a novel approach, graph representation learning via causal diffusion (CausalDiffRec) for OOD recommendation. This method enhances the model’s generalization on OOD data by eliminating environmental confounding factors and learning invariant graph representations. Specifically, we use backdoor adjustment and variational inference to infer the real environmental distribution, thereby eliminating the impact of environmental confounders. This inferred distribution is then used as prior knowledge to guide the representation learning in the reverse phase of the diffusion process to learn the invariant representation. In addition, we provide a theoretical derivation that proves optimizing the objective function of CausalDiffRec can encourage the model to learn environment-invariant graph representations, thereby achieving excellent generalization performance in recommendations under distribution shifts. Our extensive experiments validate the effectiveness of CausalDiffRec in improving the generalization of OOD data, and the average improvement is up to 10.69% on Food, 18.83% on KuaiRec, 22.41% on Yelp2018, and 11.65% on Douban datasets. Our implementation code is available at https://github.com/user683/CausalDiffRec.
Index Terms:
Graph Neural Networks, Out-of-distribution, Invariant Learning, Recommender Systems
I Introduction
Figure 1: Left and Middle: An example illustrates the popularity distribution shift, i.e., how the popularity of masks, disinfectants, exercise equipment, and electronic products changes with the COVID-19 pandemic. Right: We constructed both IID and OOD sets on the Yelp2018 dataset and compared the performance of the LightGCN model [3] on these datasets. We found a significant average performance drop (i.e., $29.03\%$) in OOD data across three metrics.
Graph Neural Networks [35, 37, 36], due to their exceptional ability to learn high-order features, have been widely applied in recommendation systems. GNN-based recommendation algorithms [1, 38, 39] learn user and item representations by aggregating information from neighboring nodes in the user-item interaction graph and then computing their similarity to predict user preferences. In addition, researchers have introduced various other techniques to continuously improve GNN-based recommendation algorithms. For example, integrating attention mechanisms [40, 8] with knowledge graphs [41] led to improving recommendation accuracy. Furthermore, the introduction of contrastive learning aims to improve the robustness of recommendation algorithms [42, 43, 42].
Despite the significant progress these methods have made in improving recommendation accuracy, most of them assume that the distribution of the test dataset and the training dataset is independently and identically distributed (IID) and focus on enhancing recommendation performance based on this assumption. Unfortunately, methods based on the above assumption fail to generalize excellent recommendation performance to out-of-distribution (OOD) data [47, 48, 49], where the distribution of the test data significantly differs from that of the training data. In Figure 1, we present a simple example to illustrate how the popularity of medical supplies changes with the environmental factor of COVID-19. Specifically, due to the pandemic, people may be required by the government to stay at home, leading to reduced outdoor activities. During this period, while purchasing medical supplies, people may also increase their demand for fitness equipment and electronic products. The recommender system might learn that users who purchase masks also frequently buy fitness equipment and electronic products, significantly increasing the popularity of these items. This correlation is driven by the common factor of the ‘pandemic’ rather than a direct causal relationship between the items. When the pandemic ends (i.e., the environmental factors change), the popularity of masks decreases, and people’s demand for fitness equipment and electronic products diminishes. The recommender system, relying on the unstable correlations learned during the training phase, may incorrectly recommend fitness equipment and electronic products to users who purchase masks, resulting in poor performance under the new distribution. Furthermore, we construct IID and OOD test sets on the Yelp2018 dataset to illustrate the changes in model performance. As shown in Figure 1, LightGCN’s [3] performance on the OOD dataset, compared to the IID dataset, shows an average decline of 29.03% across three metrics. Such issue of GNN-based models lacking robustness on OOD datasets inspires us to propose a recommendation framework with strong generalization capabilities for distribution shifts.
Several works [19, 20] have focused on improving the generalization ability of recommender systems on OOD datasets. Some researchers use causal inference to counteract shifts in the data distribution. For example, CausPref [20] is based on the NeuMF[60] method and designs invariant user preference causal learning and anti-preference negative sampling methods to improve model generalization. COR [19] uses the Variational Auto-Encoder for causal modeling by incorporating an encoder to infer unobserved user features from historical interactions. Although the aforementioned methods have improved the performance of recommendation models on out-of-distribution datasets to some extent, they are not specifically designed for GNNs, making direct migration to GNN-based methods difficult. Other researchers employ techniques such as graph contrastive learning and graph data augmentation to enhance the robustness of GNN-based recommendation algorithms, such as SGL [14], SimGCL [15], and LightGCL [31]. These methods mainly address noise or popularity bias in the data, but they struggle to achieve good recommendation performance when the test data distribution is unknown or has multiple distributions. As shown in the experimental results of Tables II, these methods perform worse than other baselines in other types of data distribution shifts. In recent literature, few GNN-based methods [23, 61] have been proposed to improve generalization when faced with multiple distributions. However, these methods lack solid theoretical support.
Given these limitations, there is an urgent need to design theoretically grounded GNN-based methods to address distribution shifts. In this paper, we use invariant learning to improve the generalization of the OOD dataset. Utilizing insights from the prior knowledge of environment distribution and invariant learning [44, 45, 46] enhances model stability across varied environments. This is achieved by acquiring invariant representations, which in turn boosts the model’s generalization capabilities and overall robustness. However, designing models based on invariant learning still faces the following two challenges:
  • 1) How to infer the distribution of underlying environments from observed user-item interaction data?
  • 2) How to recognize environment-invariant patterns amid changing user behaviors and preferences?
To address the aforementioned challenges, in this paper, we first construct the Structural Causal Model (SCM) to analyze the data generation process in recommender systems and investigate the learning process of GNN-based recommendation algorithms under data distribution shifts. We conclude that latent environmental variables can lead GNN-based algorithms to capture unstable correlations related to the environment, which is the key reason for the failure of GNN-based models to generalize on OOD data. Furthermore, we propose a novel approach called graph representation learning via Causal Diffusion (CausalDiffRec) for OOD recommendation, which leverages causal inference to eliminate unstable correlations caused by environmental variables. This approach aims to learn invariant representations across different environments, thereby achieving OOD generalization in recommender systems. Specifically, CausalDiffRec consists of three main components: an environment generator, an environment inference, and a diffusion module. The environment generator is used to create $K$ significantly different graphs to simulate data distributions under various environments. The environment inference module then infers the environment components from these generated graphs and uses them as input for the diffusion reverse stage to guide invariant graph representation learning. Finally, we provide theoretical proof that CausalDiffRec, under the conditions of invariant learning theory, can identify invariant graph representations across different environments, thereby improving generalization performance on OOD data. The contributions of this paper are concluded as follows:
  • Causal Analysis. We construct the SCM and analyze the generalization ability of GNN-based recommendation models on OOD data from the perspective of data generation. Based on our analysis and experimental results, we conclude that environmental confounders lead the model to capture unstable correlations, which is the key reason for its failure to generalize under distribution shifts.
  • Methodology. We propose a novel GNN-based method, CausalDiffRec, for OOD recommendation. CausalDiffRec primarily consists of three modules: environment generation, environment inference, and the diffusion module. The environment generation module simulates user data distributions under different conditions/environments; the environment inference module employs causal inference and variational approximation methods to infer the environment distribution; and the diffusion module is used for graph representation learning. Our theoretical analysis guarantees that optimizing the objective function of CausalDiffRec enables the model to achieve great generalization.
  • Experimental Findings. We constructed three common types of distribution shifts across four datasets and conducted comparative experiments. The experiments demonstrate that CausalDiffRec consistently outperforms baseline methods. Specifically, when dealing with OOD data, CausalDiffRec exhibits enhanced generalization capabilities, achieving a maximum metric improvement rate of 36.73% compared to the baseline methods.
II Preliminary
II-A GNN-based Recommendation
Given the observed implicit interaction matrix $\mathcal{R}\in\{0,1\}^{m\times n}$, in which $\mathcal{U}=\{u_{1},u_{2},\ldots,u_{m}\}$ represents the set of users, =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{n}\}$ represents the set of items, $m$ and $n$ denote the number of users and items, respectively. For the elements in the interaction matrix, $r_{ui}=1$ indicates an interaction between user $u$ and item $i$, otherwise 0. In GNN-based recommendation algorithms, the user-item interaction matrix $\mathcal{R}$ is first transformed into a bipartite graph $G=\{\mathcal{V},\mathcal{E}\}$. We employ $\mathcal{V}$ to represent the node set and =$\mathcal{E}=\{(u,i)|u\in\mathcal{U},i\in\mathcal{I},r_{ui}=1\}$ denotes the edge set. Given a user-item interaction graph $\mathcal{G}_{u}$ and the true user interactions $y_{u}$ with respect to (w.r.t) user $u$, the optimization objective of GNN-based methods can be expressed as:
$\mathrm{arg\;min}_{\theta}\;\mathbb{E}_{(\mathcal{G}_{u},y_{u})\sim P(G,Y)}[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})],$ (1)
where $f_{\theta}(\cdot)$ is a learner that learns representations by aggregating high-order neighbor information from the user-item interaction graph. $l$ denotes the loss function and $P(G,Y)$ represents the joint distribution of the interaction graph $G$ and true label $Y$.
II-B Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPM) [25] have been widely used in the field of image and video generation. The key idea of DDPM is to achieve the generation and reconstruction of the input data distribution through a process of gradually adding and removing noise. It leverages neural networks to learn the reverse denoising process from noise to real data distribution.
The diffusion process in recommender system models the evolution of user preferences and item information through noise addition and iterative recovery. Initially, data $\mathbf{x}_{0}$ sampled from $q(\mathbf{x})$ undergo a forward diffusion to generate noisy samples $\mathcal{\textbf{x}}_{1},\ldots,\mathbf{\textbf{x}}_{T}$ over $T$ steps. Each step adds Gaussian noise, transforming the data distribution incrementally [25]:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\right),$ (2)
where $\beta_{t}$in $(0,1)$ controls the level of the added noise at step $t$. In the reverse phase, the aim is to restore the original data by learning a model $p_{\theta}$ to approximate the reverse diffusion from $\mathbf{x}_{T}$ to $\mathbf{x}_{0}$. The process, governed by $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, uses the mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$ learned via neural networks:
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\right).$ (3)
The reverse process is optimized to minimize the variational lower bound (VLB), balancing the fidelity of reconstruction and the simplicity of the model [26]:
$\mathcal{L}_{VLB}$ $=\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\sum_{t=1}^{T}D_{KL}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))\right]$ (4)
$-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1}),$
where $D_{KL}$ denotes the Kullback-Leibler (KL) divergence. Following [25], we mitigate the training instability issue in the model by expanding and reweighting each KL divergence term in the VLB with specific parameterization. Therefore, we have the following mean squared error loss:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}_{t}}\left[\left\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}\left(\sqrt{\alpha}_{t}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon},t\right)\right\|^{2}\right],$ (5)
where $\epsilon_{t}\sim\mathcal{N}(0,\mathrm{\textbf{I}})$ is the noise for injection in forward process, $\epsilon_{\theta}(\cdot)$ denotes a function approximator that neural networks can replace, and $\alpha_{t}$ = $\prod_{t=1}^{T}1-\beta_{t}$. This framework allows the model to effectively learn noise-free representations and improve recommendation accuracy.
Figure 2: The structure causal model for GNN-based recommendation
III Methodology
In this section, we first construct SCM and identify environmental confounders as the key reason for the failure of GNN-based models to generalize on OOD (out-of-distribution) data. Subsequently, we introduce the variational inference to infer the true distribution of the environment. We use the diffusion model to learn the representation based on invariant learning. Finally, we provide rigorous theoretical proof of CausalDiffRec that can achieve great generalization. The model framework is illustrated in Figure 3.
III-A SCM of GNN-based Recommendation
To explore the reasons behind the failure of GNN-based models to generalize on OOD data, we follow previous works [27] [21] and first construct the SCM for data generation and data modeling in recommendation systems, as shown in Figure 2 (a) and Figure 2 (b). We find that environmental confounding factors are the key reason for the generalization failure of GNN-based methods. Finally, we design an intervention model in Figure 2 (c) to eliminate the impact of environmental confounding factors.
III-A1 Causal View in GNN-based Recommendation
In Figure 2, the three causal relations are derived from the definitions of data generation. The detailed causal analysis behind them is presented as follows:
  • E denotes the unobserved environmental factors (i.e., sudden hot events or policies). G and Y denote the user-item interaction graph and the true label, respectively. I is the invariant attribute of users and items unaffected by environmental factors, such as user gender and item category information. Previous work [20] has indicated that leveraging these invariant features can effectively enhance the model’s generalization capability in OOD environments.
  • E $\rightarrow$ G. The direct influence of the environment on user-item interactions can be defined as $P(G|E)$. For example, if the environmental variable is the weather, users might interact more frequently with warm clothing in a cold environment.
  • G $\rightarrow$ Y. This represents the influence of the user-item interaction graph G on the user behavior label Y. The GNN-based recommendation model $Y=f_{\theta}(G)$ defines the relation. When the model parameters $\theta$ are fixed, the mapping between G and Y is also deterministic.
  • I $\rightarrow$ Y. This indicates that the invariant user-item attributes directly influence the user behavior label $Y$. For example, a user may consistently prefer to dine at a particular restaurant, and attributes such as the restaurant’s location and name generally do not change.
  • E $\rightarrow$ Y. This denotes that the environment directly influences the user behavior label $Y$, independent of user interactions between users and items. For example, in a specific holiday environment, users may be more inclined to purchase holiday-related items, regardless of whether they have interacted with them.
In real-world scenarios, training data is collected from heterogeneous environments. Therefore, the environment directly influences the distribution of the data and the prediction result, which can be explicitly represented as $P(Y,G|E)=P(G|E)P(Y|G,E)$. If we employ $D_{\text{tr}}(E)$ to represent the training data distribution for unobserved environments, the GNN-based model, when faced with OOD data, can rewrite Eq. (1) as:
$\mathrm{arg\;min}_{\theta}\;\mathbb{E}_{e\sim D_{tr}(E),(\mathcal{G}_{u},y_{u})\sim P(G,Y|E=e)}[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})|e],$ (6)
Eq. (6) shows that environment $E$ affects the data generation used for training the GNN-based recommendation model.
III-A2 Confounding Effect of E
Figure 2 (a) and Figure 2 (b) illustrate the causal relationships in data generation and model training for graph-based recommendation algorithms. $E$ acts as the confounder and directly optimizing $P(Y|G)$ leads the GNN-based recommendation model to learn the shortcut predictive relationship between $\mathcal{G}_{u}$ and $y_{u}$, which is highly correlated with the environment $E$. During the model training process, there is a tendency to use this easily captured shortcut relationship to model user preferences. However, this shortcut relationship is highly sensitive to the environment $E$. When the environment of the test set is different from that of the training set (i.e., $D_{tr}(E)\neq D_{ts}(E)$), this relationship becomes unstable and invalid. The recommendation model that excessively learns environment-sensitive relationships in the training data will struggle to accurately model user preferences when faced with out-of-distribution data during the testing phase, resulting in a decrease in recommendation accuracy.
III-A3 Intervention
Through the above analysis, we can improve the generalization ability of GNN-based recommendation models by guiding the model to uncover stable predictive relationships behind the training data, specifically those that are less sensitive to environmental changes. Thus, we can eliminate the influence of environmental confounders on model predictions. Specifically, we learn stable correlations between user item interaction $\mathcal{G}_{u}$ and ground truth $y_{u}$ by optimizing $P_{\theta}(Y|do(G))$ instead of $P_{\theta}(Y|G)$. In causal theory, the $do$-operation signifies removing the dependencies between the target variable and other variables. As shown in Figure 2 (c), by cutting off the causal relationship between the environment variables and the user interaction graph, the model no longer learns the unstable correlations between $\mathcal{G}_{u}$ and $y_{u}$. The $do$-operation simulates the generation process of the interaction graph $G$, where environmental factors do not influence the user-item interactions. This operation blocks the unstable backdoor path $G\leftarrow E\rightarrow Y$, enabling the GNN-based recommendation model to capture the desired causal relationship that remains invariant under environmental changes.
Theoretically, $P_{\theta}(Y|do(G))$ can be computed through randomized controlled trials, which involve randomly collecting new data from any possible environment to eliminate environmental bias. However, such physical interventions are challenging. For instance, in a short video recommendation setting, it is impossible to expose all short videos to a single user, and it is also impractical to control the environment of data interactions. In this paper, we achieve a statistical estimation of $P_{\theta}(Y|do(G))$ by leveraging backdoor adjustment. The derivation process is shown as follows:
$P_{\theta}(Y|do(G))$ (7)
$=\sum_{e}P_{\theta}(Y|do(G),E=e,I)P_{\theta}(E=e|do(G))P_{\theta}(I)$
$=\sum_{e}P_{\theta}(Y|G,E=e,I)P_{\theta}(E=e|do(G))P_{\theta}(I)$
$=\sum_{e}P_{\theta}(Y|G,E=e,I)P_{\theta}(E=e)P_{\theta}(I)$
$=\sum_{e}P_{\theta}(Y|G,E=e,I)P_{\theta}(E=e,I)$
$=\mathbb{E}_{e\sim D_{tr}(E)}[P_{\theta}(Y|G,E,I)],$
through the aforementioned backdoor adjustment, the influence of the environment $E$ on the generation of $G$ can be eliminated, enabling the model to learn correlations independent of the environment. However, in recommendation scenarios, environmental variables are typically unobservable or undefined, and their prior distribution $P(E=e)$ cannot be computed. Therefore, directly optimizing the Eq. (7) is challenging.
III-B Model Instantiations
III-B1 Environment Inference
This work introduces a variational inference method and proposes a variational inference-based environment instantiation mechanism. The core idea is to use variational inference to approximate the true distribution of environments and generate environment pseudo-labels as latent variables. The following tractable evidence lower bound (ELBO) can be obtained as the learning objective:
$\mathrm{log}P_{\theta}(Y|do(G))\geq\mathcal{L}_{envInf}=\mathbb{E}_{Q_{\phi}(E|G,I)}[\mathrm{log}P_{\theta}(Y|G,E,I)]$ (8)
$-D_{KL}(Q_{\phi}(E|G,I)\parallel P_{\theta}(E)),$
where $Q_{\phi}(E|G,I)$ denotes environment estimation, which draws samples from the true distribution of the environment $E$. $D_{KL}$ represents the Kullback–Leibler (KL) divergence of the volitional distribution $Q_{\phi}(E|G,I)$ and the prior distribution $P_{\theta}$(E). $P_{\theta}(Y|G,E,I)$ is the graph representation learning module that employs the user-item interaction graph and the node attributes of users and items as input to learn invariant representations. Section III-B3 will provide a detailed introduction to the graph representation learning module. We present the derivation process of Eq. (8) as follows:
Taking the logarithm on both sides of Eq. (8) and according to Jensen’s Inequality, we have:
$\mathrm{log}P_{\theta}(Y|do(G))$ (9)
$=\mathrm{log}\;\mathbb{E}_{e\sim D_{tr}(E)}[P_{\theta}(Y|G,E,I)]$
$=\mathrm{log}\sum_{e}P_{\theta}(Y|G,E=e,I)P_{\theta}(E=e,I)\frac{Q_{\phi}(E=e|G,I)}{Q_{\phi}(E=e|G,I)}$
$\geq\sum_{e}Q_{\phi}(E=e|G,I)\;$
$\quad\quad\mathrm{log}\;P_{\theta}(Y|G,E=e,I)P_{\theta}(E=e,I)\frac{1}{Q_{\phi}(E=e|G,I)}$
$=\sum_{e}[Q_{\phi}(E=e|G,I)\;\mathrm{log}\;P_{\theta}(Y|G,E=e,I)-$
$\quad\quad\mathrm{log}\;\frac{Q_{\phi}(E=e|G,I)P_{\theta}(E=e,I)}{Q_{\phi}(E=e|G,I)}]$
$=\mathbb{E}_{Q_{\phi}(E=e|G,I)}[\mathrm{log}P_{\theta}(Y|G,E=e,I)]$
$-D_{KL}(Q_{\phi}(E=e|G,I)\parallel P_{\theta}(E=e)).$
III-B2 Invariant Pattern Recognition Mechanism
Following previous invariant learning studies [28, 48], this work proposes an invariant pattern recognition mechanism to explore invariant graph representations, encouraging the model to learn invariant correlations under distribution shifts. We make the following assumption:
Figure 3: Overall framework illustration of the proposed CausalDiffRec model. It mainly consists of three parts: 1) Environment Generator. It uses a reinforcement learning-based environment generator to generate $K$ user interaction graphs with different variances to simulate user-item interaction data under different environments. 2) Causal Diffusion. This part consists of an environment variational recommendation module and a diffusion-based graph representation learning module, which learns invariant representations of user-item interactions based on invariant learning. 3) Recommendation Module. This module initializes the recommendation model with the learned user-item representations to predict user preferences.
Assumption: For a given user-item interaction graph (i.e., data distribution $D$), these interaction data are collected from $K$ different environments $E$. User behavior patterns exist independently of the environment and can be used to generalize out-of-distribution user preference prediction. There exists an optimal invariant graph representation learning $F^{*}(\cdot)$ satisfying:
  • Invariance Property. $\forall e\in D(E),P_{\theta}(Y|F^{*}(G),E=e,I)=P(Y|F^{*}(G),I)$.
  • Sufficiency Condition. $Y=F^{*}(G)+\epsilon,\epsilon\bot E$, where $\bot$ indicates statistical independence and $\epsilon$ is random noise.
The invariance property assumption indicates that a graph representation learning model exists capable of learning invariant user-item representations across different data distribution environments. The sufficiency condition assumption means that the learned invariant representations enable the model to make accurate predictions.
III-B3 Invariant Representation Learning
This section mainly consists of an environment generator, a diffusion-based graph representation learning module, and a recommendation module. Next, we will detail how they collaborate to enhance the generalizability of GNN-based models on OOD data and improve recommendation accuracy.
Environment Generator. In real-world recommendation scenarios, training datasets are collected in various environments. However, for a single user-centric interaction graph, the training dataset comes from a single environment. We need to learn environment-invariant correlations from training data originating from different environments to achieve the generalization capability of GNN-based recommendation models under distribution shifts. To circumvent this dilemma, this paper designs an environment generator $g_{\omega_{k}}(\cdot)(1\leq k\leq K)$, which takes the user’s original interaction graph $G$ as input and generates a set of $K$ interaction graphs $\{G_{i}\}_{i=1}^{K}$ to simulate training data from different environments. The optimization objective is expressed as follows:
$\mathcal{L}_{generator}=\left[\mathrm{Var}(\mathcal{L}(g_{\omega_{k}}(G)):1\leq k\leq K)\right],$ (10)
where $Var(\cdot)$ denotes the variance and $\mathcal{L}(\cdot)$ is the loss function. Following existing work [27], we modify the graph structure by adding and removing edges. Given a Boolean matrix $B_{k}$, the adjacency matrix $A$ of the graph, and its complement $A^{\prime}$, the $k$-th generated view for the original view is $A_{k}=A+B_{k}\odot(A-A^{\prime})$. Since $B_{k}$ is a discrete matrix and not differentiable, it cannot be optimized directly. To address this issue, we borrow the idea from [27] and use reinforcement learning to treat graph generation as a decision process and edge editing as actions. Specifically, for view $k$, we consider a parameter matrix $\theta_{k}=\{\theta_{nm}^{k}\}$. For the $n$-th node, the probability of exiting the edge between it and the $m$-th node is given by:
$h(a_{nm}^{k})=\frac{\exp(\theta_{nm}^{k})}{\sum^{m^{\prime}=m}_{m^{\prime}=1}\exp(\theta_{nm^{\prime}}^{k})}.$ (11)
We then sample $s$ actions $\{b_{nmt}^{k}\}_{t=1}^{s}$ from a multinomial distribution $M(h(\alpha_{n1}^{k}),\ldots,h(\alpha_{nm}^{k}))$, which give the nonzero entries in the $n$-th row of $B_{k}$. The reward function $R(G_{k})$ can be defined as the inverse loss. We can use the reinforcement algorithm to optimize the generator with the gradient:
$\nabla_{\theta_{k}}\log h_{\theta_{k}}(A_{k})R(G_{k}),$ (12)
where $\theta_{k}$ is the model parameters and $h_{\theta_{k}}(A_{k})=\prod_{n}\prod_{t=1}^{s}h(b_{nmt}^{k})$. Optimizing Eq. (10) ensures that the generated graphs have large differences.
Causal Diffusion. Given the generated interaction graph $G_{k}=(A_{k},I)$ where $A$ is the adjacency matrix, and $I$ is the feature matrix of users or items, instead of directly using $G_{k}$ as input for diffusion, we use the encoder from the Variational Graph Autoencoder (VGAE) to compress $G_{k}$ to a low-dimensional vector $\mathbf{x}^{0}_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k})$ for subsequent environment inference and graph invariant representation leanring. The encoding process is as follows:
$q_{\psi}(\mathbf{x}^{0}_{k}|A_{k},I)=N(\mathbf{x}^{0}_{k}|\mu_{k},\sigma_{k}),$ (13)
where $\mu_{k}=GCN_{\mu}(A_{k},I)$ is matrix of mean vectors and $\sigma_{k}=GCN_{\sigma}(A_{k},I)$ denotes standard deviation. The $GCN(\cdot)$ is the graph convolution network in the graph variational autoencoder. According to the reparameterization trick, $\textbf{x}^{0}_{k}$ can be calculated as follows:
$\mathbf{x}^{0}_{k}=\mu_{k}+\sigma_{k}\odot\epsilon,$ (14)
where $\epsilon\sim N(0,I)$ and $\odot$ is the element product. Latent embedding $\mathbf{x}^{0}_{k}$ will be used as the input for the environment inference module to generate environment pseudo-labels. Meanwhile, $\mathbf{x}^{0}_{k}$ will do the forward and reverse processes in the latent space to learn the user/item embeddings in DDPM. The forward process can be calculated as:
$q(\mathbf{x}^{1:T}_{k}\mid\mathbf{x}^{0}_{k})=\prod_{t=1}^{T}q(\mathbf{x}^{t}_{k}\mid\mathbf{x}^{t-1}_{k}).$ (15)
After obtaining the environment approximation variable $z_{causal}=Q(E|G_{k},I)$ according to Eq. (8), the pair of latent variables $(z_{causal},\mathbf{x}^{T}_{k})$ to learn the invariant graph representation. We approximate the inference distribution by parameterizing the probabilistic decoder through a conditional DDPM $p_{\theta}(\mathbf{x}^{t-1}_{k}|\mathbf{x}^{T}_{k},z_{causal})$. Using DDPM, the forward process is entirely deterministic except for $t=1$. We define the joint distribution of the reverse generative process as follows:
$p_{\theta}(\mathbf{x}^{0:T}_{k}\mid z_{causal})=p(\mathbf{x}^{T}_{k})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}^{t-1}_{k}\mid\mathbf{x}^{t}_{k},z_{causal}).$ (16)
$p_{\theta}(\mathbf{x}^{t-1}_{k}|\mathbf{x})=\mathcal{N}\left(\mathbf{x}^{t-1}_{k};\mu_{\theta}(\mathbf{x}^{t}_{k},z_{causal},t),\Sigma_{\theta}(\mathbf{x}^{t}_{k},z_{causal},t)\right).$ (17)
The loss function in Eq. (5) can be rewritten as:
$\mathcal{L}_{\text{Invsimple}}=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}_{t}}\left[\left\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}^{t}_{k},z_{causal},t\right)\right\|^{2}\right],$ (18)
where $\mathbf{x}^{t}_{k}=\sqrt{\alpha}_{t}\mathbf{x}^{0}_{k}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$. After obtaining the reconstructed output vector $\mathcal{R}=\mathbf{x}^{0^{\prime}}_{k}$ from DDPM, it will be used as the input for the decoder of the variational graph decoder, which then reconstructs the input graph. The entire process is illustrated as follows:
$\hat{A}_{k}=\phi(\mathcal{R}\mathcal{R}^{\top}),$ (19)
where $\phi(\cdot)$ is the activation function (the sigmoid function is used in this paper). The VGAE is optimized by the variational lower bound:
$\mathcal{L}_{VGAE}=\mathbb{E}_{q_{\psi}(\mathbf{x}^{0}_{k}|A_{k},I)}\left[\log p_{\theta}(\hat{A}_{k}|\mathbf{x}^{0}_{k})\right]$ (20)
$-D_{KL}\left(q_{\psi}(\mathbf{x}^{0}_{k}|A_{k},I)\parallel p(\mathbf{x}^{0}_{k})\right).$
Prediction and Joint Optimization. Using the well-trained diffusion model to sample the final embeddings for user preference modeling:
$\hat{r}_{u,i}=e^{\top}_{u}e_{i},$ (21)
where $e_{u}$ and $e_{i}$ denote the final user embedding w.r.t $u$-th user and item embedding w.r.t $i$-th item, respectively. Without loss of generality, LightGCN is used as the recommendation backbone, and Bayesian Personalized Ranking (BRP) loss is employed to optimize the model parameters:
$\mathcal{L}_{rec}=\sum_{u,v^{+},v^{-}}-\log\sigma(\hat{r}_{u,v^{+}}-\hat{r}_{u,v^{-}}),$ (22)
where $(u,v^{+},v^{-})$ is a triplet sample for pairwise recommendation training. $v^{+}$ represents positive samples from which the user has interacted, and $v^{-}$ are the negative samples that are randomly drawn from the set of items with which the user has not interacted, respectively. We use a joint learning strategy to optimize CausalDiffRec:
$\mathcal{L}=$ $\mathcal{L}_{rec}+\lambda_{1}\cdot\mathcal{L}_{\text{generator}}$ (23)
$+\lambda_{2}\cdot(\mathcal{L}_{\text{VGAE}}+\mathcal{L}_{\text{Invsample}})+\lambda_{3}\cdot\mathcal{L}_{\text{envInf}},$
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyper-parameters.
In section IV, we provide the theoretical proof that our proposed CausalDiffRec can get rid of the unstable correlation and learn OOD generalized representations when satisfying the Assumption. The pseudocode for the model training process is shown in Algorithm 1.
IV Theoretical Analysis
CausalDiffRec aims to learn the optimal generator $F^{*}(\cdot)$ as stated in the assumption in section III-B2, thereby obtaining invariant graph representations to achieve OOD generalization in recommendation performance under data distribution shifts. Before starting the theoretical derivation, let’s do some preliminary work. For the convenience of theoretical proof, we rewrite Eq. (23) as:
$\mathrm{argmin}_{\theta}\;(\mathcal{L}_{\text{task}}+\mathcal{L}_{\text{infer}}),$ (24)
where $\mathcal{L}_{task}=\mathcal{L}_{rec}+\mathcal{L}_{\text{generator}}+\mathcal{L}_{\text{VGAE}}+\mathcal{L}_{\text{Invsample}}$ and $\mathcal{L}_{\text{infer}}=\mathcal{L}_{\text{envInf}}$, and for the derivation convenience, we temporarily ignore the penalty coefficient. The $\mathcal{L}_{\text{task}}$ and $\mathcal{L}_{\text{infer}}$ can be further abstracted as:
$\mathcal{L}_{\text{task}}=\mathrm{arg\;min}_{\theta}\mathbb{E}_{e\sim D_{tr}(E),(\mathcal{G}_{u},y_{u})\sim P(Y,G|E=e)}$ (25)
$\quad[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})]$
$\mathcal{L}_{\text{infer}}=\mathrm{min}_{q(Y|z_{causal})}\mathrm{Var}\{\mathbb{E}_{e\sim D_{tr}(E),(\mathcal{G}_{u},y_{u})\sim P(Y,G|E=e)}$
$\quad[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})|do(\mathcal{G}_{u})]\}.$
We follow the proof technique from [28] and show the optimality of the Eq. (24) with the following two propositions that can achieve OOD recommendation.
Proposition 1
Minimizing Eq. (24) promotes the model’s adherence to the Invariance Property and the Sufficient Condition outlined in Assumption (in Sec. III-B2).
Proposition 2
Optimizing Eq. (24) corresponds to minimizing the upper bound of the OOD generalization error described in Eq. (6).
Proposition 1 and Proposition 2, respectively, avoid strong hypotheses and ensure that the OOD generalization error bound of the learned model is within the expected range. In fact, this can also be explained from the perspective of the SCM model in Figure 2. Optimizing Eq. (6) eliminates the negative impact of unstable correlations learned by the model, which are caused by latent environments, on modeling user preferences. At the same time, it enhances the model’s ability to learn invariant causal features across different latent environments. Proofs for Proposition 1 and Proposition 2 are shown as follows:
Before starting the proof, we directly follow previous work [28, 27, 50, 51, 52] to propose the following lemma, using information theory to interpret the invariance property and sufficient condition in Assumption and to assist in the proof of Proposition 1.
Using the Mutual Information $\mathbb{I}(;)$, the invariance property and sufficient condition in Assumption can be equivalently expressed as follow lemma:
Lemma 1
(1) Invariance: $\forall e\in D(E),P_{\theta}(Y|\mathcal{P}^{*}_{Inv},E=e,I)=P(Y|\mathcal{P}^{*}_{Inv},I)\Leftrightarrow\mathbb{I}(Y;E|\mathcal{P}^{*}_{Inv},I)=0$ where $\mathcal{P}^{*}_{Inv}=F^{*}(G)$. (2) Sufficiency: $\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)$ is maxmized.
For the invariance property, it is easy to get the following equation:
$\mathbb{I}(Y;E|\mathcal{P}^{*}_{\text{Inv}},I)$ (26)
$=\mathbb{E}_{\mathcal{P}^{*}_{\text{Inv}},I}\left[\mathbb{D}_{\text{KL}}\left(P(Y,E|\mathcal{P}^{*}_{\text{Inv}},I)\|P(Y|\mathcal{P}^{*}_{\text{Inv}},I)P(E|\mathcal{P}^{*}_{\text{Inv}},I)\right)\right]$
For the sufficient condition, we use the method of contradiction and prove it through the following two steps:
First, we prove that for $Y$, $\mathcal{P}^{*}_{Inv}$, and $I$ satisfying $\mathcal{P}^{*}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, they also satisfy that $\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)$ is maximized. We use the method of contradiction to prove this. Assume $\mathcal{P}^{*}_{Inv}\neq\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, and there exists $\mathcal{P}^{\prime}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, where $\mathcal{P}^{\prime}_{Inv}\neq\mathcal{P}^{*}_{Inv}$. We can always find a mapping function $M$ such that $\mathcal{P}^{\prime}_{Inv}=M(\mathcal{P}^{*}_{Inv},R)$, where $R$ is a random variable. Then we have:
$\mathbb{I}(Y;\mathcal{P}^{\prime}_{Inv},I)$ $=\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},R,I)$ (27)
$=\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)+\mathbb{I}(Y;R|\mathcal{P}^{*}_{Inv},I).$
Since $R$ is a random variable and does not contain any information about $Y$, we have $\mathbb{I}(Y;R|\mathcal{P}^{*}_{Inv},I)=0$. Therefore:
$\mathbb{I}(Y;\mathcal{P}^{\prime}_{Inv},I)=\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I).$ (28)
This leads to a contradiction.
Next, we prove that for $Y$, $\mathcal{P}^{*}_{Inv}$, and $I$ satisfying $\mathcal{P}^{*}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, they also satisfy that $\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)$ is maximized. Assume $\mathcal{P}^{*}_{Inv}\neq\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, and there exists $\mathcal{P}^{\prime}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$, where $\mathcal{P}^{\prime}_{Inv}\neq\mathcal{P}^{*}_{Inv}$. We have the following inequality:
$\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)\leq\mathbb{I}(Y;\mathcal{P}^{\prime}_{Inv},I).$ (29)
From this, we can deduce that:
$\mathcal{P}^{\prime}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I),$ (30)
where contradicts $\mathcal{P}^{*}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$. Since the assumption leads to a contradiction, and the assumption does not hold. Therefore, $\mathcal{P}^{*}_{Inv}=\arg\max_{\mathcal{P}_{Inv}}\mathbb{I}(Y;\mathcal{P}_{Inv},I)$ holds. This proves that $\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)$ is maximized. The lemma 1 is complicated proven.
Proof of Proposition 1. First, optimizing the first term $\mathcal{L}_{\text{task}}$ in Eq. (24) enables the model to satisfy the sufficient condition. Analyzing the SCM in Figure 2(c), we have the fact that $\mathrm{max}_{q(z|G,I)}\mathbb{I}(Y,z_{causal})$ is equivalent to $\mathrm{min}_{q(z|G,I)}\mathbb{I}(Y,G|Z_{causal})$, as we use $do(G)$ to eliminate the unstable correlations between Y and G caused by the latent environment. We have:
$\mathbb{I}(Y,G|z_{causal})$ $=D_{KL}(p(Y|G,E)\|p(Y|z_{causal},E))$ (31)
$=D_{KL}(p(Y|G,E)\|p(Y|z_{causal}))$
$-D_{KL}(p(Y|z_{causal},E)\|q(Y|z_{causal}))$
$\leq D_{KL}(p(Y|G,E)\|p(Y|z_{causal})).$
Based on the above derivation, we have:
$\mathbb{I}(Y,G|z_{causal})\leq$ (32)
$\mathrm{min}_{q(Y|z_{causal})}\;D_{KL}(p(Y|G,E)\|p(Y|z_{causal})).$
Besides, we have:
$D_{KL}(p(Y|G,E)\|p(Y|z_{causal}))$ (33)
$=\mathbb{E}_{e\in D_{tr}(E)}\mathbb{E}_{(G,Y)\sim p(G,Y|e)}$
$\quad\mathbb{E}_{z_{causal}\sim q(z_{causal}|G,I)}\left[\mathrm{log}\frac{q(Y|G,e)}{p(Y|z_{causal})}\right]$
$\leq\mathbb{E}_{e\in D_{tr}(E)}\mathbb{E}_{(G,Y)\sim p(G,Y|e)}$
$\left[\mathrm{log}\frac{p(Y|G,e)}{\mathbb{E}_{z_{causal}\sim q(z_{causal}|G,I)}q(Y|z_{causal})}\right]\mathrm{(Jensen\;Inequality)}.$
Finally, we reach:
$\mathrm{min}_{q(Y|z_{causal})}\;D_{KL}(p(Y|G,E)\|p(Y|z_{causal}))$ (34)
$\Leftrightarrow\mathrm{arg\;min}_{\theta}\mathbb{E}_{e\sim D_{tr}(E),(\mathcal{G}_{u},y_{u})\sim P(Y,G|E=e)}[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})].$
Thus, we have demonstrated that minimizing the expectation term ($\mathcal{L}_{task}$) in Eq. (24) is equivalent to minimizing the upper bound of $\mathbb{I}(Y;G\mid z_{causal})$. This results in maximizing $\mathbb{I}(Y;\mathcal{P}^{*}_{Inv},I)$, thereby helping to ensure that the model satisfies the Sufficient Condition.
Next, we prove that optimizing the first term $\mathcal{L}_{task}$ in Eq. (24) enables the model to satisfy the Invariance Property. Similar to Eq. (31), we have:
$\mathbb{I}(Y;E=e\mid z_{causal})$ (35)
$=D_{KL}(p(Y\mid z_{causal},e)\parallel p(Y\mid z_{causal}))$
$=D_{KL}(p(Y\mid z_{causal},E)\parallel\mathbb{E}_{e\in D(E)}[p(Y\mid z_{causal},e)])$
$=D_{KL}(q(Y\mid z_{causal})\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})])$
$-D_{KL}(q(Y\mid z_{causal})\parallel p(Y\mid z_{causal},e))$
$-D_{KL}(\mathbb{E}_{e\in D(E)}[p(Y\mid z_{causal},e)]\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})])$
$\leq D_{KL}(q(Y\mid z_{causal})\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})]).$
Besides, the last term in Eq. (35) can be further expressed as:
$D_{KL}(q(Y\mid z_{causal})\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})])$ (36)
$=\mathbb{E}_{e\in D_{tr}(E)}\mathbb{E}_{(G,Y)\sim p(G,Y|e)}\mathbb{E}_{z_{causal}\sim q(z_{causal}|G,I)}$
$\quad\left[\mathrm{log}\frac{p(Y|G,e)}{\mathbb{E}_{e\in D(E)}q(Y|z_{causal})}\right]\mathrm{(Jensen\;Inequality)}$
$\leq\mathbb{E}_{e\in D(E)}[|l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})-\mathbb{E}_{e\in D(E)}[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})]|]$
where the last term in Eq. (36) is the upper bound for the $D_{KL}(q(Y\mid z_{causal})\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})])$. Finally, we have:
$\mathrm{min}_{q(Y|z_{causal})}D_{KL}(q(Y\mid z_{causal})\parallel\mathbb{E}_{e\in D(E)}[q(Y\mid z_{causal})])$ (37)
$\Leftrightarrow\mathrm{min}_{q(Y|z_{causal})}\mathrm{Var}\{\mathbb{E}_{e\sim D_{tr}(E),(\mathcal{G}_{u},y_{u})\sim P(Y,G|E=e)}$
$[l(f_{\theta}(\mathcal{G}_{u};\theta),y_{u})|do(\mathcal{G}_{u})]\}.$
Hence, minimizing the variance term ($\mathcal{L}_{risk}$) in Eq. (24) effectively reduces the upper bound of $\mathbb{I}(Y;E=e\mid z_{causal})$. Thereby ensuring the model adheres to the Invariance Property.
Proof of Proposition 2. Optimizing Eq. (24) is tantamount to reducing the upper bound of the OOD generalization error in Eq. (6). Let $q(Y\mid G)$ represent the inferred variational distribution of the true distribution $p(Y\mid G,E)$. The OOD generalization error can be quantified by the KL divergence between these two distributions:
$D_{KL}(p(Y|G,E)\|q(Y|G))$ (38)
$=\mathbb{E}_{e\in D_{tr}(E)}\mathbb{E}_{(Y,G)\sim p(Y,G|E)}\mathrm{log}\frac{p(Y|G,E=e)}{q(Y|G)}.$
Following previous work, we use information theory to assist in the proof of Proposition 2. We propose the lemma 2 to rewrite the OOD generalization, which is shown as follows:
Lemma 2
The out-of-distribution generalization error is limited by:
$D_{KL}(p(Y|G,E)\|q(Y|G))\leq D_{KL}[p(Y|G,E)\|q(Y|z_{causal})],$ (39)
where $q(Y|z_{causal})$ is the inferred variational environment distribution. The proof of Lemma 2 is shown as:
$D_{KL}(p(Y|G,E=e)\|q(Y|z_{causal}))$ (40)
$=\mathbb{E}_{e\in D(E)}\mathbb{E}_{(Y,G)\sim p(G,Y|E=e)}\left[\mathrm{log}\frac{p(Y|G,E=e)}{q(Y|G)}\right]$
$=\mathbb{E}_{e\in D(E)}\mathbb{E}_{(Y,G)\sim p(G,Y|E=e)}$
$\quad\left[\mathrm{log}\frac{p(Y|G,E=e)}{\mathbb{E}_{z_{causal}\sim q(z_{causal}|G,I)}q(Y|z_{causal})}\right]$
$\leq\mathbb{E}_{e\in D(E)}\mathbb{E}_{(Y,G)\sim p(G,Y|E=e)}$
$\quad\mathbb{E}_{z_{causal}\sim q(z_{causal}|G,I)}\mathrm{log}\frac{p(Y|G,e)}{q(Y|z_{causal})}$
$=D_{KL}[p(Y|G,E)\|q(Y|z_{causal})],$
The Lemma 2 has been fully proven. Based on Lemma 1 and Proposition 1, the Eq. (24) can be adapted as:
$\mathrm{min}_{q(z_{causal}|G,I),q(Y,z_{causal})}$ (41)
$D_{KL}(p(Y|G,E=e)\|q(Y|z_{causal}))+\mathbb{I}(Y,E=e|z_{causal})$
Hence, according to Lemma 2, we confirm that minimizing Eq. (24) is tantamount to minimizing the upper bound of the OOD generalization error in Eq. (6), meaning that:
$\mathrm{argmin}_{\theta}(\mathcal{L}_{task}+\mathcal{L}_{infer})\Leftrightarrow\mathrm{min}_{q(z_{causal}|G,I),q(Y,z_{causal})}$ (42)
$\quad D_{KL}(p(Y|G,E=e)\|q(Y|z_{causal}))$
$+\mathbb{I}(Y,E=e|z_{causal})\;(\mathbb{I}(Y,E=e|z_{causal})\;is\;non-negative)$
$\geq\mathrm{min}_{q(z_{causal}|G,I),q(Y,z_{causal})}$
$\quad D_{KL}(p(Y|G,E=e)\|q(Y|z_{causal}))$
$\geq D_{KL}(p(Y|G,E)\|q(Y|G)).$
The Proposition 2 is completely proven.
V Experiments
In this section, we conducted extensive experiments to validate the performance of CausalDiffRec and address the following key research questions:
  • RQ1: How does CausalDiffRec compare to the state-of-the-art strategies in both OOD and IID test evaluations?
  • RQ2: Are the proposed components of CausalDiffRec effective for OOD generalization?
  • RQ3: How do hyperparameter settings affect the performance of CausalDiffRec?
Algorithm 1 Training of CausalDiffRec under Multiple Environments
1:  Input: The user-item interaction graph $G(\mathcal{V},\mathcal{E})$ and node feature matrix $\mathcal{X}$; Using the $\omega$, $\theta_{1}$, and $\theta_{2}$ to initial environment generator $g_{\omega}(\cdot)$, environment $P_{\theta_{1}}(\cdot)$, and graph representation learner (ie., sampling approximator) $f_{\theta_{2}}(\cdot)$, respectively.
2:  while not converged do
3:     for all $u\in U$ do
4:        for all $k\in\{1,2,\ldots,K\}$ do
5:           Get the modified modified graphs $G_{k}$ by Eq. (10);
6:           Infer the causal environment label $z_{causal}$ from Eq. (9);
7:           Obtain the damage representation $\textbf{x}^{t}_{K}$ by Eq. (15);
8:           // Forward process
9:           Sample $\textbf{x}^{t-1}_{k}$ by feeding $z_{causal}$ and $\textbf{x}^{t}_{k}$ into $f_{\theta_{2}}(z_{causal},\textbf{x}^{t}_{k})$;
10:           // Reverse process
11:           Calculate $f_{\theta_{2}}(\textbf{x}^{t-1}_{k})$ via Eq. (18) ;
12:           Calculate the gradients w.r.t. the loss in Eq. (27);
13:        end for
14:     end for
15:     Average the gradients over $|U|$ users and $K$ environments;
16:     Update $\omega$, $\theta_{1}$, and $\theta_{2}$ via AdamW optimizer;
17:  end while
18:  Output: $g_{\omega}(\cdot)$, $P_{\theta_{1}}(\cdot)$, and $f_{\theta_{2}}(\cdot)$.
V-A Experimental Settings
Datasets. We evaluate our proposed CausalDiffRec on four real-world datasets: Food [^1] https://www.aclweb.org/anthology/D19-1613/, KuaiRec [^2] https://kuairec.com Yelp2018333https://www.yelp.com/dataset, and Douban [^4] https://www.kaggle.com/datasets/ comprise raw data from the Douban system. Detailed statistics of the datasets are presented in Table 1. More data processing details can be seen in the code implementation link. Following exiting work [24], we constructed OOD test sets for three common scenarios of distribution shift:
  • Temporal shift. We sort the dataset in descending order by timestamp and used each user’s most recent 20% of interactions as the OOD test set. The first 60% of the data based on interaction time is used for the training set. Food is used for this type of shift.
  • Exposure shift. In KuaiRec, the small matrix is fully exposed and used for the OOD test set, and the big matrix is collected from the online platform and used for the training dataset, resulting in a distributional shift.
  • Popularity shift. We randomly select 20% of the interactions to create the OOD test set, ensuring that item popularity follows a uniform distribution. The training set maintains a long-tail distribution. This type of shift was applied to the Yelp2018 and Douban datasets.
TABLE I: Detailed statistics for each dataset.
Dataset #Users #Items #Interactions Density
Food 7,809 6,309 216,407 $4.4\times 10^{-3}$
KuaiRec 7,175 10,611 1,153,797 $1.5\times 10^{-3}$
Yelp2018 8,090 13,878 398,216 $3.5\times 10^{-3}$
Douban 8,735 13,143 354,933 $3.1\times 10^{-3}$
Baselines. We compare the CusalDiffRec with the state-of-the-art models: LightGCN[14], SGL [14], SimGCL[15], LightGCL[31], InvPref [22], InvCF [33], AdvDrop [23], AdvInfo [34], and DRO [24].
Hyperparameter Settings. We implement our CausalDiffRec in Pytorch. All experiments are conducted on a single RTX-4090 with 24G memory. Following the default hyperparameter search settings of the baselines, we expand their hyperparameter search space and tune the hyperparameters. For our CausalDiffRec, we tune the learning rates in $\{1e-3,1e-4,1e-5\}$. The number of diffusion steps varies between 10 and 1000, and the diffusion embedding size is chosen in $\{8,16,32,64\}$. Additional hyperparameter details are available in our released code.
TABLE II: The performance comparison between the baselines and CausalDiffRec on the four datasets with three data distribution shifts. The best results are highlighted in bold, and the second-best results are underlined. ’Impro.’ denotes the relative improvements of CausalDiffRec over the second-best results.
Dataset Metric LightGCN SGL SimGCL LightGCL InvPref InvCF CDR AdvDrop AdvInfo DRO Ours Impro.
Food R@10 0.0234 0.0198 0.0233 0.0108 0.0029 0.0382 0.0260 0.0240 0.0227 0.0266 0.0251 1.99%
N@10 0.0182 0.0159 0.0186 0.0101 0.0014 0.0237 0.0195 0.0251 0.0135 0.0205 0.0296 24.89%
R@20 0.0404 0.0324 0.0414 0.0181 0.0294 0.0392 0.0412 0.0371 0.0268 0.0436 0.0464 6.03%
N@20 0.0242 0.0201 0.0249 0.0121 0.0115 0.0240 0.0254 0.0237 0.0159 0.0279 0.0306 9.68%
KuaiRec R@10 0.0742 0.0700 0.0763 0.0630 0.0231 0.1023 0.0570 0.1014 0.1044 0.0808 0.1116 6.90%
N@10 0.5096 0.4923 0.5180 0.4334 0.2151 0.2242 0.2630 0.3290 0.4302 0.5326 0.6474 21.25%
R@20 0.1120 0.1100 0.1196 0.1134 0.0478 0.1034 0.0860 0.1214 0.1254 0.1266 0.1631 28.83%
N@20 0.4268 0.4181 0.4446 0.4090 0.2056 0.2193 0.2240 0.3289 0.4305 0.4556 0.5392 18.35%
Yelp2018 R@10 0.0014 0.0027 0.0049 0.0022 0.0049 0.0004 0.0011 0.0027 0.0047 0.0044 0.0067 36.73%
N@10 0.0008 0.0017 0.0028 0.0015 0.0030 0.0026 0.0006 0.0017 0.0024 0.0029 0.0039 30.00%
R@20 0.0035 0.0051 0.0106 0.0054 0.0108 0.0013 0.0016 0.0049 0.0083 0.0076 0.0120 11.65%
N@20 0.0016 0.0026 0.0047 0.0026 0.0049 0.0008 0.0008 0.0024 0.0038 0.0041 0.0055 11.24%
Douban R@10 0.0028 0.0022 0.0086 0.0070 0.0052 0.0030 0.0014 0.0051 0.0076 0.0028 0.0094 9.30%
N@10 0.0015 0.0013 0.0045 0.0038 0.0026 0.0012 0.0007 0.0021 0.0042 0.0011 0.0050 11.11%
R@20 0.0049 0.0047 0.0167 0.0113 0.0093 0.0033 0.0200 0.0046 0.0103 0.0038 0.0197 17.96%
N@20 0.0019 0.0020 0.0073 0.0050 0.0038 0.0013 0.0019 0.0021 0.0053 0.0015 0.0079 8.22%
V-B Overall Performance (RQ1)
This section compares CausalDiffRec’s performance and baselines under various data shifts and conducts a performance analysis.
Evaluation on temporal shift: Table II shows that CausalDiffRec significantly outperforms SOTA models on the Food dataset, with improvements of 1.99%, 24.89%, 6.03%, and 9.68% in Recall and NDCG. This indicates CausalDiffRec’s effectiveness in handling temporal shift. DRO also excels in this area, with a 15% improvement over LightGCN in NDCG@20, due to its robust optimization across various data distributions. CDR surpasses GNN-based models thanks to its temporal VAE-based architecture, capturing preference shifts from temporal changes.
Evaluation on exposure shift: In real-world scenarios, only a small subset of items is exposed to users, leading to non-random missing interaction records. Using the fully exposed KuaiRec dataset, CausalDiffRec consistently outperforms baselines, with improvements ranging from 6.90% to 28.83%, indicating its capability to handle exposure bias. DRO and AdvInfoNce also show superior performance in NDCG and Recall metrics, enhancing the generalization of GNN-based models and demonstrating robustness compared to LightGCN.
Evaluation on popularity shift. We compare model performance on the Yelp2018 and Douban datasets, showing that our model significantly outperforms the baselines. On Douban, CausalDiffRec achieves 8.22% to 17.96% improvement, and on Yelp2018, the improvements range from 11.24% to 36.73%. Methods using contrastive learning (e.g., SimGCL, LightGCL, AdvInfoNce) outperform other baselines in handling popularity shifts. This is because the InfoNCE loss helps the model learn a more uniform representation distribution, reducing bias towards popular items. InvPref performs best among the baselines on Yelp2018, using clustering for contextual labels, unlike our variational inference approach. Our method, tailored for graph data, aggregates neighbor information for better recommendation performance than matrix factorization-based methods.
Additionally, in Table III, we report the performance of CausalDiffRec compared to several baseline models that use LightGCN as the backbone. From the table, we observe the following: 1) These baseline models outperform LightGCN on IID datasets; 2) CausalDiffRec outperforms all baseline models across all metrics. This indicates that CausalDiffRec also performs well on IID datasets. We attribute the performance improvement to our use of data augmentation and the incorporation of auxiliary information in modeling user preferences.
In summary, the analysis of experimental results demonstrates that our proposed CausalDiffRec can handle different types of distribution shifts and achieve good generalization.
TABLE III: Model performance comparison on IID datasets.
Dataset Ablation R@10 R@20 N@10 N@20
KuaiRec LightGCN 0.0154 0.0174 0.0272 0.0210
AdvDrop 0.0431 0.0258 0.0469 0.0276
AdvInfo 0.0514 0.0298 0.0518 0.0300
DRO 0.0307 0.0226 0.0505 0.0291
CausalDiffRec 0.0567 0.0472 0.0634 0.0707
Yelp2018 LightGCN 0.0023 0.0022 0.0039 0.0029
AdvDrop 0.0061 0.0051 0.0063 0.0050
AdvInfo 0.0052 0.0040 0.0062 0.0049
DRO 0.0069 0.0550 0.0086 0.0065
CausalDiffRec 0.0102 0.0120 0.0182 0.0134
TABLE IV: Outcomes from ablation studies on four datasets. The top-performing results are highlighted in bold, while those that are second-best are underlined.
Dataset Ablation R@10 R@20 N@10 N@20
Food LightGCN 0.0234 0.0404 0.0182 0.0242
w/o Gen. 0.0165 0.0259 0.0114 0.0148
w/o Env. 0.0084 0.0144 0.0077 0.0098
CausalDiffRec 0.0251 0.0409 0.0296 0.0306
KuaiRec LightGCN 0.0808 0.1266 0.5326 0.4556
w/o Gen. 0.0966 0.1571 0.0445 0.3078
w/o Env. 0.0047 0.1740 0.0697 0.0784
CausalDiffRec 0.1116 0.1631 0.0674 0.5392
Yelp2018 LightGCN 0.0014 0.0035 0.0008 0.0016
w/o Gen. 0.0041 0.0054 0.0037 0.0042
w/o Env. 0.0027 0.0058 0.0042 0.0043
CausalDiffRec 0.0067 0.0120 0.0039 0.0055
Douban LightGCN 0.0028 0.0049 0.0015 0.0019
w/o Gen. 0.0044 0.0079 0.0030 0.0045
w/o Env. 0.0044 0.0070 0.0023 0.0031
CausalDiffRec 0.0094 0.0197 0.0050 0.0079
V-C In-depth Analysis (RQ2)
In this section, we conduct ablation experiments to study the impact of each component of CausalDiffRec on recommendation performance. The main components include the environment generator module and the environment inference module. Additionally, we use t-SNE to visualize the item representations captured by the baseline model and CausalDiffRec, to compare the models’ generalization capabilities on OOD data.
Ablation studies. Table IV presents the results of the ablation study that compares LightGCN, CausalDiffRec, and its two variants: ’w/o Gen.’ (without the environment generator) and ’w/o Env.’ (without the environment inference). The results show that removing these modules causes a significant drop in all metrics across four datasets. For example, on Yelp2018, Recall, and NDCG decreased by 148.15% and 126.83%, respectively, demonstrating the effectiveness of CausalDiffRec based on invariant learning theory for enhancing recommendation performance on OOD datasets. Additionally, even without the modules, CausalDiffRec still outperforms LightGCN on popularity shift datasets (Yelp2018 and Douban) due to the effectiveness of data augmentation and environment inference. However, on the Food and KuaiRec datasets, removing either module results in worse performance than LightGCN, likely due to multiple biases in these datasets. Without one module, the model struggles to handle multiple data distributions, leading to a performance drop. Overall, the ablation experiments highlight the importance of all modules in CausalDiffRec for improving recommendation performance and generalizing on OOD data.
Visualization analysis. In Figure 4 and Figure 5, we used t-SNE to visualize the item representations learned by LightGCN, SimGCL, and CausalDiffRec on Douban and Yelp2018 datasets to better observe our model’s ability to handle distribution shifts. Following previous work [24], we recorded the popularity of each item in the training set and designated the top 10% most popular items as ’popular items’ and the bottom 10% as ’unpopular items’. It is obvious that the embeddings of popular and unpopular items learned by LightGCN still exhibit a gap in the representation space. In contrast, the embeddings learned by CausalDiffRec are more evenly distributed within the same space. This indicates that CausalDiffRec can mitigate the popularity shift caused by popular items. Additionally, we found that the embeddings of popular and unpopular items learned by SimGCL are more evenly distributed compared to LightGCN. This is because contrastive learning can learn a uniform representation distribution.
Figure 4: Visualization of user embedding distributions using various methods on the Douban dataset. CausalDiffRec ensures that hot items and cold items have representations that are nearly co-located within the same space.
Figure 5: Visualization of user embedding distributions using various methods on the Yelp2018 dataset. CausalDiffRec ensures that hot items and cold items have representations that are nearly co-located within the same space.
      (a) Food
      (b) KuaiRec
      (c) Yelp2018
Figure 6: Effects of the number of diffusion steps T.
       (a) $\lambda_{1}$ on Food
      (b) $\lambda_{2}$ on Food
       (c) $\lambda_{3}$ on Food
       (d) $\lambda_{1}$ on KuaiRec
      (e) $\lambda_{2}$ on KuaiRec
      (f) $\lambda_{3}$ on KuaiRec
Figure 7: Effect of the coefficients in the objective loss on Food and KuaiRec dataset.
      (a) Food
      (b) KuaiRec
      (c) Yelp2018
      (d) Food
      (e) KuaiRec
      (f) Yelp2018
Figure 8: Effects of the number of environments K.
V-D Hyperparameter Investigation (RQ3)
Effect of Diffusion Step T. We conducted experiments to investigate the impact of the number of diffusion steps on performance; in CausalDiffRec, we used the same number of steps in the forward and reverse phases. we compare the performance with $T$ changing from 10 to 500. We present the results in Figure 6, and we have the following findings:
  • When the number of steps is chosen within the range {10, 50, 100}, CausalDiffRec achieves the best performance across all datasets. We find that appropriately increasing the number of steps significantly improves Recall@20 and NDCG@20. These performance enhancements are mainly attributed to the diffusion enriching the representation capabilities of users and items.
  • Nevertheless, as we continue to increase the number of steps, the model will face overfitting issues. For example, on the food and yelp2018 datasets, Recall@20 and NDCG@20 consistently decrease. Although there is an upward trend on KuaiRec, the optimal solution is not achieved. Additionally, it is evident that more steps also lead to longer training times. We should carefully adjust the number of steps to find the optimal balance between enhancing representation ability and avoiding overfitting.
Effect of Penalty Coefficient. Figure 7 reports the experimental results of varying the coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ on food and KuaiRec. We compare the performances with all coefficients changing from 1e-1 to 1e-4. Further, we have the following conclusions:
  • $\lambda_{1}\in[1e-1,1e-2]$ typically strikes a good balance and delivers excellent performance. When further reducing the value of $\lambda_{1}$, the model’s performance declined. We speculate that this led to overfitting. Similarly, we found that larger values of $\lambda_{3}$ can also achieve better recommendation performance. Therefore, the value range of $\lambda_{3}$ is the same as that of $\lambda_{1}$.
  • $\lambda_{3}$ directly regularizes the capability of the graph representation learning module. We found that for different datasets, values that are too large or too small do not allow the model to learn optimal representations. Taking everything into consideration, we set $\lambda_{3}$ to 1e-3 to achieve good generalization.
Effect of the number of Environments. Figure 8 shows the impact of the number of environments on the model’s performance. We can see that as the number of environments increases, the performance of CausalDiffRec improves across the three datasets. This indicates that more environments help enhance the model’s generalization on OOD (Out-of-Distribution) data. However, as the number of environments further increases, performance declines, which we believe is due to the model overfitting to too many environments.
VI Related Work
VI-A GNN-based Recommendation
Recent developments in graph-based recommender systems have leveraged graph neural networks to model user-item interactions as a bipartite graph, enhancing recommendation accuracy through complex interaction capture [1, 2, 3]. Notably, LightGCN focuses on neighborhood aggregation without additional transformations, while other approaches employ attention mechanisms to prioritize influential interactions [4, 5, 6, 7, 8, 9]. Further research explores non-Euclidean spaces like hyperbolic space to better represent user-item relationships [10, 11]. Knowledge graphs also enhance these systems by integrating rich semantic and relational data directly into the recommendation process [13, 12]. Despite these advancements, graph-based systems often struggle with out-of-distribution data due to the IID assumption and are challenged by multiple distribution shifts [14, 15, 16, 17, 18]. Additionally, contrastive learning methods in these systems rely on a fixed paradigm that lacks robust theoretical support, limiting adaptability to varied data shifts. However, the aforementioned models are trained on datasets where the training and test data distributions are drawn from the same distribution, leading to generalization failure when facing OOD data.
VI-B Diffusion based Recoomendation
The integration of diffusion processes into recommender systems leverages diffusion mechanisms to model dynamic propagation of user preferences and item information through interaction networks, enhancing recommendation accuracy and timeliness [53, 54, 55, 56, 57, 58, 59]. These models capture evolving user behaviors and have shown potential in various recommendation contexts, from sequential recommendations to location-based services. For instance, DiffRec [53] applies diffusion directly for recommendations, while Diff-POI [57] models location preferences. Furthermore, approaches like DiffKG [58] and RecDiff [59] utilize diffusion for denoising entity representations in knowledge graphs and user data in social recommendations, respectively, enhancing the robustness and reliability of the systems. These studies underscore diffusion’s suitability for advanced representation learning in recommender systems. However, these methods cannot solve the OOD problem.
VI-C Out-of-Distribution Recommendation
Researchers have focused on recommendation algorithms for out-of-distribution (OOD) data. COR [19] infers latent environmental factors in OOD data. CausPref [20] learns invariant user preferences and causal structures using anti-preference negative sampling. CaseQ [21] employs backdoor adjustment and variational inference for sequential recommendations. InvPref [22] separates invariant and variant preferences by identifying heterogeneous environments. However, these methods don’t directly apply to graph-based recommendation models and fail to address OOD in graph structures. AdaDrop [23] uses adversarial learning and graph neural networks to enhance performance by decoupling user preferences. DRO [24] integrates Distributionally Robust Optimization into Graph Neural Networks to handle distribution shifts in graph-based recommender systems. Distinct from these GNN-based methods, this paper explores how to use the theory of invariant learning to design GNN-based methods with good generalization capabilities.
VII Conclusion
In this paper, we propose a novel GNN-based model for OOD recommendation, called CausalDiffRec, designed to learn environment-invariant graph representations to enhance the generalization of recommendation models on OOD data. CausalDiffRec employs the backdoor criterion from causal inference and variational inference methods to eliminate the influence of environmental confounders. It uses a diffusion-based sampling strategy to learn graph representations. Our approach is grounded in invariant learning theory, and we provide theoretical proof that by optimizing the objective function in CausalDiffRec, the model is encouraged to identify invariant graph representations across environments, thereby improving its generalization performance on OOD data. Experiments conducted on four real-world datasets demonstrate that the proposed CausalDiffRec framework outperforms baseline models. In addition, ablation studies further validate the effectiveness of the model.